Circular No. 8482

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN NGC 4945 C. Jacques and E. Pimentel, Belo Horizonte, Brazil, report their discov-

ery of an apparent supernova (mag 12.8) on unfiltered CCD images taken on Feb. 8.22 and 10.23 UT with a 0.30-m Schmidt-Cassegrain reflector in the course of the CEAMIG/REA Supernovae Search. The new object is located at α = 13^{h}04^{m}44^{s}.06, δ = −49^{o}33′59′′.8 (equinox 2000.0), which is 407′′ west and 351′′ south of the center of NGC 4945, but which also lies in a rather rich area of foreground Milky Way stars in Centaurus. Nothing in visible on a CCD image taken by Jacques on 2004 June 20.93 (limiting mag 18.5) or on a red Digitized Sky Survey plate from 1976.

SUPERNOVAE 2005Q, 2005S, 2005Y, 2005Z, 2005ad N. Morrell, M. Hamuy, and G. Folatelli, Carnegie Supernova Project;

and F. Olivares, University of Chile, report that spectra (range 380–930 nm) were obtained on Feb. 4 and 8 UT with the Las Campanas 2.5-m du Pont Telescope (+ WFCCD spectrograph), showing that SNe 2005Q, 2005Y, 2005Z, and 2005ad (cf. IAUC 8473, 8476, 8479) are all of type II. SNe 2005ad, 2005Z, and 2005Q display blue continua with strong P-Cyg profiles in the Balmer lines and He I 587.6-nm, indicative of young events. SN 2005ad seems to be very young, based on the very blue continuum. SN 2005Y shows the Balmer lines and the Fe II 516.9-nm absorption, implying a more evolved event. The expansion velocities, derived from the minimum of the Hβ absorption (assuming for the parent galaxies the recession velocities listed in the NED database) are: 2005Q, 11500 km/s; 2005Y, 6800; 2005Z, 11600; 2005ad, 7400. A spectrum of SN 2005S reveals it to be a type-Ia supernova, 2–3 weeks after maximum light.

SUPERNOVAE 2005R, 2005X, AND 2005Y R. J. Foley and A. V. Filippenko, University of California, Berkeley,

report that inspection of CCD spectra (range 500–1000 nm), obtained on Feb. 11 UT with the Keck II 10-m telescope (+ DEIMOS), shows that SN 2005R (IAUC 8473) is of type IIn; the broad Hα emission line has a full-width-at-half-maximum of ∼ 2400 km/s. SN 2005X (IAUC 8476) is of type Ia, ∼ 2 weeks past maximum brightness; the redshift of its host galaxy is measured to be z = 0.0745. SN 2005Y (IAUC 8476) is of type II, showing a prominent Hα P-Cyg profile. After removing the host-galaxy recession velocity of 4900 km/s (from NED), the minimum of the broad Hα absorption is measured to be −7400 km/s relative to its rest wavelength.

2005 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green